OpenStudy (anonymous):

how do you solve this ?? help sqrt 1-(1/2)^2

OpenStudy (anonymous):



OpenStudy (anonymous):

Simplify (1/2)^2

OpenStudy (anonymous):

1/4

OpenStudy (anonymous):

Now convert 1 to a fraction to have the same denominator as 1/4

OpenStudy (anonymous):

4/4

OpenStudy (anonymous):

Now subtract

OpenStudy (anonymous):

3/4

OpenStudy (anonymous):

Okay, so we're left with: \[\frac{ \sqrt{3} }{ \sqrt{4} }\] What's the final answer?

OpenStudy (anonymous):